8. (8 pts) For a solution that contains 197 mg of a nonelectrolyte and 168 mL of water. The solution froze at-1.29℃ and Kr-1.8601m, what is the molar mass of the nonelectrolyte?
Answer #1

∆Tf = i × Kb × m

m = molality = mol/ mass of solvent (Kg)

m = mol/0.168

Mass of solvent (water) = v × d

= 168×1= 168 g = 0.168 kg

0 - (-1.29) = 1×1.86× mole/0.168

Mole = 0.116 mol

Molar mass = mass/ mole

= 197×10-3 g/.116

= 1.69 g/mol